### v2

# Emanuel Castelo Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Oscar Defrain Aix-Marseille Université, CNRS, LIS, 13009 Marseille, France Guilherme C. M. Gomes LIRMM, Université de Montpellier, CNRS, Montpellier, France Universidade Federal de Minas Gerais, Belo Horizonte, Brazil

Abstract
Enumerating minimal dominating sets with polynomial delay in bipartite graphs is a long-standing open problem. To date, even the subcase of chordal bipartite graphs is open, with the best known algorithm due to Golovach, Heggernes, Kanté, Kratsch, Sæther, and Villanger running in incremental-polynomial time. We improve on this result by providing a polynomial delay and space algorithm enumerating minimal dominating sets in chordal bipartite graphs. Additionally, we show that the total and connected variants admit polynomial and incremental-polynomial delay algorithms, respectively, within the same class. This provides an alternative proof of a result by Golovach et al. for total dominating sets, and answers an open question for the connected variant. Finally, we give evidence that the techniques used in this paper cannot be generalized to bipartite graphs for (total) minimal dominating sets, unless $\mathsf{P=NP}$, and show that enumerating minimal connected dominating sets in bipartite graphs is harder than enumerating minimal transversals in general hypergraphs.
Keywords: algorithmic enumeration, minimal dominating sets, connected dominating sets, total dominating sets, chordal bipartite graphs, sequential method, polynomial delay.
1 Introduction
Hypergraphs are encountered in countless areas of computer science or discrete mathematics. Formally, they consist of a pair $\mathcal{H}=(V(\mathcal{H}),E(\mathcal{H}))$ where $V(\mathcal{H})$ is a set of elements called vertices, and $E(\mathcal{H})$ is a family of subsets of $V(\mathcal{H})$ called hyperedges (or simply edges). Hypergraphs are known as graphs when their edges are of size precisely two. A transversal in a hypergraph $\mathcal{H}$ is a set $T\subseteq V(\mathcal{H})$ which intersects every edge of $\mathcal{H}$. It is called minimal if it is inclusion-wise minimal, i.e., if none of its proper subsets is a transversal. The problem of enumerating the minimal transversals of a hypergraph, usually denoted by Trans-Enum, is one of the most important problems in algorithmic enumeration, for it has far-reaching implications in various areas of computer science such as logic, database theory, data profiling, or artificial intelligence [KPS93, EG95, GMKT97, BFL+22]; we refer the reader to [EG95] and the surveys [EMG08, GDVL17] for a comprehensive overview on the fundamental and practical aspects of this problem.
In a typical enumeration problem such as Trans-Enum, a first observation to be made is that the number of solutions can be exponential in both $n$, the number of vertices, and $m$, the number of edges. Thus, asking for running times which are polynomial in $n+m$ is meaningless. Rather, we are interested in algorithms performing in output-polynomial time, i.e., that run in a time which is polynomial in both the input size, and the output size. More strict notions of tractability can be required, such as incremental-polynomial times, and polynomial delay. We say that an algorithm runs in incremental-polynomial time if it produces the $i^{\text{th}}$ solution in a time which is polynomial in the input size plus $i$. An algorithm is said to run with polynomial delay if the times before the first output, after the last output, and in between two consecutive outputs are bounded by a polynomial in the input size only. To date, the best-known algorithm solving Trans-Enum is due to Fredman and Khachiyan [FK96] and runs in quasi-polynomial incremental time. Since then, output-polynomial time algorithms were obtained for several important classes, including those generalizing that of bounded degree [EGM03], or bounded edge size [KBEG07].
Dominating sets are among the most studied objects in graphs. A dominating set in a graph $G$ is a set $D\subseteq V(G)$ such that every vertex is either in $D$, or adjacent to a vertex of $D$. Again, it is called minimal if it is inclusion-wise minimal. The problem of enumerating minimal dominating sets in graphs is denoted by Dom-Enum. Note that being a dominating set of $G$ can be rephrased as being a transversal of each closed neighborhood of $G$, so, Dom-Enum is a special case of Trans-Enum with a linear number of hyperedges.
The problem Dom-Enum has gained lots of interest since it has been shown, perhaps surprisingly, to be polynomially equivalent to Trans-Enum, even when restricted to cobipartite graphs [KLMN14]. In particular, this work has initiated a fruitful line of research of characterizing which graph classes admit output-polynomial algorithms, in the hope of better understanding the underlying structure that makes Trans-Enum tractable. For example, polynomial-delay algorithms were obtained for various graph classes including chordal graphs [KLM+15a] (with linear delay on some subclasses [KLMN14, DN19]), graphs of bounded degeneracy [EGM03], line graphs [KLM+15b], graphs of bounded clique-width [Cou09], or graph classes related to permutation graphs and their generalizations [KLMN12, GHK+18, BDMN20]. Incremental-polynomial time algorithms were obtained for graphs of bounded conformality [KBEG07], $C_{\leq 6}$-free graphs [GHKV15], chordal bipartite graphs [GHK+16a], $(C_{4},C_{5},\text{claw})$-free graphs [KLMN12], or graphs arising in geometrical contexts [ERR19, GHK+18] including unit-disks and unit-square graphs. Output-polynomial time algorithms are known for $\log n$-degenerate graphs [EGM03], or graphs of bounded clique number [BDH+20].
In [KLMN14], the authors also include two natural variants of domination in their study: total and connected dominating sets. A total dominating set is a set $D$ such that every vertex of $G$ has a neighbor in $D$. In other words, elements of $D$ cannot “dominate” themselves and need a neighbor in the set. A connected dominating set is a dominating set that induces a connected subgraph. Note that any connected dominating set is a total dominating set, as long as it has more than one element, making the intersection of these two notions irrelevant to study. Perhaps not surprisingly, the authors in [KLMN14] show that the problems of enumerating minimal total and connected dominating sets, denoted TDom-Enum and CDom-Enum, respectively, are equivalent to Trans-Enum when restricted to split graphs. These equivalences on a relatively restricted graph class suggest that these variants of Dom-Enum are challenging, and may explain why they have received less attention in the literature.
Regarding total dominating sets, a first observation is that they are transversals of open neighborhoods. This implies an incremental quasi-polynomial time algorithm for the problem using the results of Fredman and Khachiyan [FK96]. However, few output-polynomial time algorithms are known for specific instances of TDom-Enum. Among them, we cite the class of chordal bipartite graphs, which admits a polynomial-delay algorithm for the problem [GHK+16a]
[…]
Note that, in the literature, particular attention has been put into characterizing the status of these problems on bipartite, split, and cobipartite graphs. This is because such instances contain the incidence graph of any hypergraph, which is fundamentally challenging and makes interesting links with Trans-Enum. However, and in contrast to the split and cobipartite cases, the case of bipartite graphs has been particularly challenging with the question of the (in)tractability of these problems raised in various works [KN14, GHK+16a]. Only recently an output-polynomial time algorithm has been provided for Dom-Enum in a generalization of bipartite graphs [BDH+20]. Still the problem is not known to admit a polynomial-delay algorithm. Concerning TDom-Enum and CDom-Enum, it is also open, to the best of our knowledge, whether they can be solved with polynomial delay, or even in output-polynomial time, in bipartite graphs. Among natural subclasses of bipartite graphs, the case of chordal bipartite graphs has been studied as it admits nice structural properties analogous to chordal graphs, such as perfect elimination orderings and linearly many minimal separators. We note, however, that even in this particular class, the existence of a polynomial-delay algorithm for Dom-Enum seems open, and has been left open in [GHK+16a].
The following two natural questions arise from the literature.
Question 1.
Can the minimal dominating sets of a chordal bipartite graph be enumerated with polynomial delay?
Question 2.
Can the minimal connected dominating sets of a chordal bipartite graph be enumerated in output-polynomial time?
Only the case of minimal total dominating sets has been settled for chordal bipartite graphs, with the following result.
{restatable}
[[GHK+16a]]theoremthmtds The minimal total dominating sets of a chordal bipartite graph be enumerated with polynomial delay.
Our results and techniques. This paper is devoted to providing a positive answer to Questions 1 and 2, almost completing the picture for chordal bipartite graphs. Collaterally, we provide an alternative proof of Question 2 involving small modifications of our algorithm for minimal dominating sets; we note that the obtained algorithm has similar time bounds as the one in [GHK+16a].
Our algorithms for Dom-Enum (Theorem 2) and TDom-Enum (Question 2) are based on the ordered generation framework (also known as sequential method in various works) which has been introduced in [EGM03] as a generalization of the algorithm in [TIAS77] (see also [JYP88] for a concise description). This approach has since then proved to be fruitful in various contexts related to domination and minimal transversals [CKMU19, BDH+20, DNV21]. Roughly, the idea is to decompose the instance with respect to an ordering of its elements, and, given the partial solutions of the instance induced by the $i$ first elements, to extend them to the next element of the instance. To achieve polynomial delay, a parent-child relation on partial solutions is defined. Then, the resulting tree is traversed in a DFS manner, and only the leaves (which correspond to actual solutions) are output. Overall, the technique allows to reduce the general enumeration to the enumeration of the children of a partial solution, and is particularly suitable for instances admitting good elimination orderings. In our case, we use the weak-elimination ordering of chordal bipartite graphs (see Section 2) to solve children generation in polynomial time, yielding a polynomial-delay algorithm for Dom-Enum and TDom-Enum in that class.
Concerning CDom-Enum, we show that the minimal separators of a chordal bipartite graph define a hypergraph of bounded conformality (Lemma 5), a parameter introduced by Berge [Ber84] and that has been successfully used in the design of tractable algorithms for domination [KLMN12] and minimal transversals enumeration [KBEG07]. This, combined with the fact that chordal bipartite graphs have polynomially many minimal separators, which can be listed in polynomial time (see Section 2), allow us to obtain an incremental-polynomial time algorithm for CDom-Enum in that class (Theorem 4).
Additionally, we give evidence that the sequential method cannot be directly used to generalize our results to bipartite graphs. More specifically, we show that the problem of generating the children of a given partial solution is intractable for Dom-Enum (Lemma 6) and TDom-Enum (Lemma 7). For CDom-Enum we make a stronger statement (Theorem 5): the problem is at least as hard as Trans-Enum, which was not known to the best of our knowledge, despite the reduction being straightforward.
Organization of the paper. We introduce preliminary notions and properties dealing with chordal bipartite graphs and minimal separators in Section 2. We describe the sequential approach in Section 3, and provide our algorithms for Dom-Enum and TDom-Enum in Section 4. The arguments for an incremental-polynomial time algorithm for CDom-Enum in chordal bipartite graphs are given in Section 5. Limitations of the sequential method in bipartite graphs are discussed in Section 6, and the reduction from Trans-Enum to CDom-Enum in bipartite graphs is given in Section 7. We discuss open questions in Section 8.
2 Preliminary notions
We assume the reader to be familiar with standard terminology from graph and hypergraph theory, and refer to the two monographs [Die12, Ber84] for notations not included below.
Graphs. All graphs considered in this paper are simple and finite. If $v$ is a vertex of $G$ then we denote by $N(v)$ its open neighborhood, by $N[v]$ its closed neighborhood, and define its distance-2 neighborhood as $N^{2}(v):=N[N[v]]\setminus N[v]$. The first two notions are naturally extended to subsets by $N[S]:=\bigcup_{v\in S}N[S]$ and $N(S):=(\bigcup_{v\in S}N[S])\setminus S$. An independent set is a set of pairwise non-adjacent vertices. A graph is bipartite if its vertex set $V(G)$ can be partitioned into two independent sets $X,Y$. It is called bipartite chain if, in addition, there exist orderings $x_{1},\dots,x_{p}$ and $y_{1},\dots,y_{q}$ of $X$ and $Y$, respectively, such that $N(x_{i})\subseteq N(x_{j})$ for all $1\leq i\leq j\leq p$, and $N(y_{i})\supseteq N(y_{j})$ for all $1\leq i\leq j\leq q$.
Hypergraphs. Let $\mathcal{H}$ be a hypergraph. Given a vertex $v\in V(\mathcal{H})$, the set of edges that contain $v$, also called edges incident to $v$, is denoted by $\operatorname{{\sf inc}}(v)$. A transversal of $\mathcal{H}$ is a subset $T\subseteq V(\mathcal{H})$ such that $E\cap T\neq\emptyset$ for all $E\in E(\mathcal{H})$, i.e., every edge is incident to (or hit by) a vertex of $T$. The hypergraph induced by a set $X\subseteq V(\mathcal{H})$ of vertices is defines as $\mathcal{H}[X]=(X,\{E\in E(\mathcal{H}):E\subseteq X\})$. It is minimal if it is minimal by inclusion, or equivalently, if $T\setminus\{v\}$ is not a transversal for any $v\in T$. The set of minimal transversals of $\mathcal{H}$ is denoted by $Tr(\mathcal{H})$. The private edges of an element $v$ in a transversal $T$ are those edges of $\mathcal{H}$ which are only hit by $v$ in $T$; we denote the set of such edges by $\operatorname{{\sf priv}}(T,v)$. It is well-known (see e.g., [Ber84]) that a set $T\subseteq V(\mathcal{H})$ is a minimal transversal of $\mathcal{H}$ if and only if $T$ is a transversal and every $v$ in $T$ has a private edge. We formally state the problem of enumerating minimal transversals of a hypergraph as follows.
{problemgen}
Note that in the problem above, $\mathcal{H}$ can be considered inclusion-wise minimal (or Sperner) as this does not change its set of minimal transversals [EMG08].
Domination. We recall the different notions of domination given in the introduction. A dominating set (or DS for short) in a graph $G$ is a set $D\subseteq V(G)$ such that $V(G)=N[D]$. It is called connected if the graph $G[D]$ is connected. A total dominating set in $G$ is a set $D\subseteq V(G)$ such that $V(G)=\bigcup_{v\in D}N(v)$. A (resp. connected, total) dominating set is said to be minimal if it is inclusion-wise minimal. We denote by $\operatorname{MDS}(G)$ (resp. $\operatorname{MCDS}(G)$, $\operatorname{MTDS}(G)$) the families of minimal such sets in a graph $G$.
In this paper, we are interested in the following enumeration problems.
{problemgen}
{problemgen}
{problemgen}
We recall that the precise complexity status of each of the problems above is widely open, the first one admitting a quasi-polynomial time algorithm [FK96, KLMN14], while the third is conjectured to be intractable [BBHK15]. Their study has been initiated in [KLMN14], where it is noted that each of the problems is a particular instance of Trans-Enum: using the family of (closed and open) neighborhoods for the first two, and a more involved hypergraph for the third one, as formalized in Propositions 13. See also [Ber84].
Proposition 1.
For every graph $G$, $\operatorname{MDS}(G)=Tr(\{N[v]:v\in V(G)\})$.
Proposition 2.
For every graph $G$, $\operatorname{MTDS}(G)=Tr(\{N(v):v\in V(G)\})$.
Minimal separators. A separator in a graph $G$ is a subset $S\subseteq V(G)$ such that $G-S$ is disconnected. In the following, let $\mathcal{S}(G):=\{S\subseteq V:G-S~{}\text{is not connected and}~{}\forall S^{\prime}\subset S,G-S^{\prime}~{}\text{is connected}\}$ denote the set of (inclusion-wise) minimal separators of $G$. Then we have the following.
Proposition 3 ([KLMN14]).
For every graph $G$, $\operatorname{MCDS}(G)=Tr(\mathcal{S}(G))$.
Minimal separators are not to be confused with the slightly different notion of minimal $a$-$b$ separators which are minimal sets $S$ such that $a$ and $b$ lie in different components of $G-S$. Indeed, while every minimal separator is a minimal $a$-$b$ separator, the converse is not always true: minimal $a$-$b$ separators may contain another $c$-$d$ separator as a subset; see [Gol04] for an early mention of this observation, and [BDK+24] for a thorough discussion on this matter. A simple example is the cycle on four vertices with a pendant edge [BDK+24]. Moreover, while enumerating $a$-$b$ minimal separators can be done with polynomial delay [SL97, KK98, BBC00], enumerating minimal separators is a hard task [BDK+24]. However, and as is noted in [KLMN14], CDom-Enum reduces to Trans-Enum for graph classes admitting polynomially many $a$-$b$ separators. [^1] It should be noted that the statement [KLMN14, Corollary 33] presents an inaccuracy where by minimal separators it should be understood minimal $a$-$b$ separators. Indeed, listing minimal separators is hard [BDK+24] while enumerating minimal $a$-$b$ separators is tractable [KK98]. The present inaccuracy comes from an ambiguity and common confusion in the literature around this terminology [SL97, KLP11]. Let $\mathcal{G}$ be the hypergraph where each hyperedge is an $a$-$b$ minimal separator of $G$; using the above-cited algorithms for enumerating $a$-$b$ minimal separators, we can construct $\mathcal{G}$ in polynomial time on $|\mathcal{G}|+|V(G)|+|E(G)|$; moreover, note that $\mathcal{S}(G)=\operatorname{Min}_{\subseteq}\mathcal{G}$. Indeed, recall that $Tr(\mathcal{H})=Tr(\mathcal{G})$ for any pair of hypergraphs $\mathcal{H},\mathcal{G}$ such that for any $E\in E(\mathcal{G})$ there exists $E^{\prime}\in E(\mathcal{H})$ with $E^{\prime}\subseteq E$. In other words, $Tr(\mathcal{H})=Tr(\operatorname{Min}_{\subseteq}\mathcal{H})$ for any hypergraph $\mathcal{H}$ [Ber84].
Chordal bipartite graphs. A bipartite graph is chordal bipartite if it does not contain induced cycles of length greater than four [Gol04]. Note that chordal bipartite graphs are not chordal, as they may contain the induced cycle on four vertices. This terminology should be regarded as an analogue to bipartite graphs of chordality, as bicliques are to cliques.
A vertex $v$ in a graph $G$ is called weak-simplicial if $N(v)$ is an independent set, and for any two $x,y\in N(v)$, we have $N(x)\subseteq N(y)$ or $N(y)\subseteq N(x)$; two such vertices $x,y$ are said to be comparable. A weak-simplicial elimination ordering of $G$ is an ordering $v_{1},\dots,v_{n}$ of its vertices such that, for all $1\leq i\leq n$, $v_{i}$ is weak-simplicial in $G_{i}=G\left[\{v_{i},\dots,v_{n}\}\right]$. The following characterization is due to [KWUA19] and may be regarded as a simplification of an earlier characterization and recognition algorithm by Uehara [Ueh02].
Proposition 4 ([Ueh02, KWUA19]).
A graph is chordal bipartite if and only if it admits a weak-simplicial elimination ordering, which can be computed in polynomial time.
Additionally, from the above definition it is easily seen that if $G$ is chordal bipartite, and $v$ is a weak-simplicial vertex, then the graph $G[N^{2}(v)]$ induced by the distance-2 neighborhood of $v$ is bipartite chain; see [KWUA19, Lemma 3]. A characterization based on separators is given in [GG78], where it is proved that a bipartite graph is chordal bipartite if and only if every minimal edge separator induces a complete bipartite subgraph. In [KLP11], the following analogue of the forward implication is given for minimal $a$-$b$ separators (hence for minimal separators).
Proposition 5 ([KLP11]).
If $G$ is chordal bipartite and $S$ is an $a$-$b$ minimal separator then $G[S]$ is complete bipartite.
In the same paper, the authors show that chordal bipartite graphs have at most $O(n+m)$ minimal $a$-$b$ separators [KLP11]. Since they can be listed with polynomial delay [KK98], we derive a polynomial time procedure listing $a$-$b$ minimal separators and filtering out the non-minimal ones. We formalize this in the next statement that will be of use in Section 5.
Proposition 6.
Let $G$ be an $n$-vertex chordal bipartite graph. Then its set $\mathcal{S}(G)$ of minimal separators can be enumerated in polynomial time in $n$.
Finally, we will need two extra properties of minimal separators in general and in chordal bipartite graphs. Let $S$ be a separator of $G$. We say that $C$ is close to $S$ if every vertex in $S$ has at least one neighbor in $C$. It is folklore that minimal $a$-$b$ separators $S$ are characterized by the existence of two close components in $G-S$; see [SL97, Gol04, KLP11] or [KK94] for a proof. We derive one implication for minimal separators since, in particular, they are $a$-$b$ minimal separators for some specific pairs $a,b$ of vertices. Note however that the converse is not true.
Proposition 7.
Let $G$ be a graph and $S$ be a minimal separator of $G$. Then there exist two components $C_{1},C_{2}$ in $G-S$ that are close to $S$.
Finally, the following will be of use in Section 5. We state it for minimal separators but this more generally holds for minimal $a$-$b$ separators [KLP11].
Proposition 8 ([KLP11]).
Let $G$ be chordal bipartite with bipartition $(A,B)$. Let $S$ be a minimal separator of $G$ and $C$ be a component of $G-S$ that is close to $S$. If $S\cap A\neq\emptyset$ then there exists a vertex $x$ in $C$ with $N(x)\cap S=S\cap A$.
3 The sequential method
We describe the sequential method. As by Propositions 13, Dom-Enum, TDom-Enum, and CDom-Enum can be seen as particular instances of Trans-Enum we choose to follow the presentation of [BDM24], and refer the reader to [EGM03, EHR08, BDH+20] for more details on this approach.
Let $\mathcal{H}\subseteq 2^{V}$ be a hypergraph on vertex set $V=\{v_{1},\dots,v_{n}\}$. For $1\leq i\leq n$ we denote by $V_{i}:=\{v_{1},\dots,v_{i}\}$ the set of its first $i$ vertices, by $\mathcal{H}_{i}:=\mathcal{H}[V_{i}]$ the hypergraph induced by $V_{i}$, and set $V_{0}:=\mathcal{H}_{0}:=Tr(H_{0}):=\emptyset$. In addition, for $S\subseteq V_{i}$ and $v\in S$, we note $\operatorname{{\sf priv}}_{i}(S,v)$ the set of private edges of $v$ with respect to $S$ in $\mathcal{H}_{i}$. Given a minimal transversal $T\in Tr(\mathcal{H}_{i+1})$ for $i<n$, we denote by $\operatorname{{\sf parent}}(T,i+1)$ the parent of $T$ with respect to $i+1$ obtained after repeatedly removing from $T$ the vertex of smallest index $v$ such that $\operatorname{{\sf priv}}_{i}(T,v)=\emptyset$. By definition, every $T$ has a unique parent. Lastly, we denote by $\operatorname{{\sf children}}(T^{\star},i)$, for $T^{\star}\in Tr(\mathcal{H}_{i})$, the family of $T\in Tr(\mathcal{H}_{i+1})$ such that $T^{\star}=\operatorname{{\sf parent}}(T,i+1)$.
The next two lemmas are central to the description of the algorithm: they define a tree structure where nodes are pairs $(T,i)$ with $T\in Tr(\mathcal{H}_{i})$ and that will be traversed in a DFS manner in order to produce every leaf $T\in Tr(\mathcal{H})=Tr(\mathcal{H}_{n})$ as a solution. We refer to [BDM24] for the proofs of these statements, which follow from the definition of the $\operatorname{{\sf parent}}$ relation.
Lemma 1.
Let $10\leq i\leq n-$ 1 and $T\in Tr(\mathcal{H}_{i+1})$, then $\operatorname{{\sf parent}}(T,i+1)\in Tr(\mathcal{H}_{i})$.
Lemma 2.
Let $10\leq i\leq n-$ 1 and $T^{\star}\in Tr(\mathcal{H}_{i})$, then either:
  • $T^{\star}\in Tr(\mathcal{H}_{i+1})$ and consequently $\operatorname{{\sf parent}}(T^{\star},i+1)=T^{\star}$; or
  • $T^{\star}\cup\{v_{i+1}\}\in Tr(\mathcal{H}_{i+1})$ and $\operatorname{{\sf parent}}(T^{\star}\cup\{v_{i+1}\},i+1)=T^{\star}$.
Finally, the whole framework culminates in the following statement that reduces Trans-Enum to the enumeration of the children of a given pair $(T,i)$ with $T\in Tr(\mathcal{H}_{i})$, $i<n$, while preserving polynomial delay and space.
Theorem 1.
Let $f,s:\mathbb{N}^{2}\rightarrow\mathbb{Z}_{+}$ be two functions. If there is an algorithm that, given a hypergraph $\mathcal{H}$ on vertex set $V=\{v_{1},\dots,v_{n}\}$ and $m$ edges, an integer $i\in\{1,\dots,n-1\}$, and $T^{\star}\in Tr(\mathcal{H}_{i})$, enumerates $\operatorname{{\sf children}}(T^{\star},i)$ with delay $f(n,m)$ and using $s(n,m)$ space, then there is an algorithm that enumerates $Tr(\mathcal{H})$ with $\mathcal{O}(nf(n,m)$ delay and $\mathcal{O}(ns(n,m))$ space.
In the following, given $T^{\star}$ and $i$ as above, let
$\Delta_{i+1}:=\{E\in\mathcal{H}_{i+1}:E\cap T^{\star}=\emptyset\}\subseteq\operatorname{{\sf inc}}_{i+1}(v_{i+1})$
be the family of hyperedges of $\mathcal{H}_{i+1}$ that are not hit by $T^{\star}$. The following shows that children may be found by considering minimal extensions to $\mathcal{H}_{i+1}$ of minimal transversals of $\mathcal{H}_{i}$.
Lemma 3.
If $T\in\operatorname{{\sf children}}(T^{\star},i)$ then $T=T^{\star}\cup X$ for some $X\in Tr(\Delta_{i+1})$.
Proof.
Let $T\in\operatorname{{\sf children}}(T^{\star},i)$. Then $T^{\star}$ is obtained from $T$ by greedily removing elements that have no private edge in $\mathcal{H}_{i}$. Let $X$ be the set of removed elements. Note that since $T^{\star}$ is transversal of $\mathcal{H}_{i}$, $X$ may only have private edges among those $E\in\mathcal{H}_{i+1}$ that are not intersected by $T^{\star}$, i.e., among $\Delta_{i+1}$. Then, either $\Delta_{i+1}=\emptyset$ in which case $T=T^{\star}$ by Lemma 2, or $X$ is non empty and intersects each edge in $\Delta_{i+1}$. By minimality of $T$, we derive that $X$ is minimal with this property, as desired. ∎
As a corollary, in order to enumerate $\operatorname{{\sf children}}(T^{\star},i)$ for some $i\in\{1,\dots,n-1\}$ and $T^{\star}\in Tr(\mathcal{H}_{i})$, one can simply enumerate all the minimal transversals of $\Delta_{i+1}$ and filter out those that do not provide children of $T^{\star}$. In the following, we call minimal extensions of $(T^{\star},i)$ such sets. Note, however that the number of minimal transversal of $\Delta_{i+1}$ may be greater (typically exponential) than the number of actual children, hence that this approach fails to efficiently produce children in general. Moreover, this approach is intractable in general, with instances admitting non-trivial children if and only if NP-complete problems admit a positive answer; see [BDM24] or Section 6. We nevertheless have the following which says that the approach is still tractable if we can bound the number of minimal extension by a polynomial in the input size, and which is in essence what is done in [EGM03].
Corollary 1.
If there is an algorithm that, given $T^{\star}\in Tr(\mathcal{H}_{i})$, $i<n$, enumerates all $X\in Tr(\Delta_{i+1})$ in time and space which are polynomial in $n+m$, then there is one that lists $Tr(\mathcal{H})$ with polynomial delay and space.
Note that, if $\mathcal{H}$ is $d$-degenerate and $v_{1},\dots,v_{n}$ is an optimal degeneracy ordering, then minimal extensions have size at most $d$. In [EGM03], the authors use this fact to provide a polynomial delay algorithm that lists minimal transversals in $O(1)$-degenerate hypergraphs relying on Corollary 1. See [EGM03, BDM24] for the definition of these notions. In this paper, we combine this corollary together with the elimination ordering of chordal bipartite graphs provided by Proposition 4 to show that Dom-Enum and TDom-Enum can be solved with polynomial delay as well in this class.
4 Minimal extension for (total) dominating sets
In this section, we apply the sequential method to Trans-Enum restricted to hypergraphs of closed neighborhoods of chordal bipartite graphs to show that Question 1 admits a positive answer. More specifically, we use the weak-simplicial elimination ordering of chordal bipartite graphs to restrict the instance of minimal extensions enumeration to a variant of domination in bipartite chain graphs, which is tractable. By Corollary 1, this allows us to obtain Theorem 2.
In the following, let $G$ be a chordal bipartite graph on $n$ vertices and $v_{1},\dots,v_{n}$ be a weak-simplicial ordering of its vertices as provided by Proposition 4. Let $\mathcal{H}$ be its hypergraph of closed neighborhoods, $i\in\{1,\dots,n-1\}$ be an integer and $T$ be a minimal transversal of $\mathcal{H}_{i}$ as in the statement of Corollary 1. We will show how to enumerate the minimal transversals of $\Delta_{i+1}$ in polynomial time and space, to derive the desired result.
First note that since the edges of $\Delta_{i+1}$ belong to $\mathcal{H}_{i+1}$ and are incident to $v_{i+1}$, they may be equal to $N[v_{i+1}]$ if $N[v_{i+1}]\subseteq V_{i+1}$, or to $N[u]$ for some neighbor $u$ of $v$ if $N[u]\subseteq V_{i+1}$. From now on, let $B$ be the set of vertices $u\in N(v_{i+1})$ such that $N[u]\in\Delta_{i+1}$, and let
$R:=\left(\bigcup\Delta_{i+1}\right)\setminus B$
be the remaining vertices that lie in some edge of $\Delta_{i+1}$. Consequently $B\subseteq N(v_{i+1})$ and $R\subseteq N(v_{i+1})\cup N^{2}(v_{i+1})$. Then, by Proposition 4, the graph $H=G_{i}[R\cup B]$ is bipartite chain. We denote by $(X,Y)$ its bipartition where $B\subseteq X\subseteq N(v_{i+1})$, and by $x_{1},\dots,x_{p}$ and $y_{1},\dots,y_{q}$ the respective orderings of $X$ and $Y$ satisfying $N(x_{j})\cap Y\subseteq N(x_{\ell})\cap Y$ and $N(y_{\ell})\cap X\subseteq N(y_{j})\cap X$ for all $j<\ell$; see Section 2 and Figure 1 for an illustration of the situation.
Figure 1: The bipartite chain instance induced by $X$ and $Y$ with $v_{i+1}$ depicted as well. Blue vertices are those whose closed neighborhoods are incident to $v_{i+1}$ and included in $V_{i+1}$ as a subset. Red vertices in $X$ have neighbors out of $V_{i}$.
In what remains we shall call blue the vertices in $B$, red those in $R$, and characterize minimal transversals of $\Delta_{i+1}$ as particular sets analogous to red-blue dominating sets where red and blue vertices may be selected in order to dominate blue vertices.
Proposition 9.
Let $T\in Tr(\Delta_{i+1})$. Then
  • $|T\cap R\cap N(v_{i+1})|\leq 1$; and
  • $|T\cap N^{2}(v_{i+1})|\leq 1$.
Proof.
First, note that if $x$ belongs to $T\cap R\cap N(v_{i+1})$ then it is incident to at most one edge in $\Delta_{i+1}$ (as $H$ is bipartite), namely $N[v_{i+1}]$, in which case $N[v_{i+1}]$ is a private edge of $x$. Hence, no two distinct vertices $x,y$ may lie in $T\cap R\cap N(v_{i+1})$ by minimality of $T$, yielding the first inequality.
Now, recall that $N^{2}(v_{i+1})\subseteq Y$ hence that one of $N(y)\cap X\subseteq N(y^{\prime})\cap X$ or $N(y)\cap X\supseteq N(y^{\prime})\cap X$ holds for any pair $y,y^{\prime}\in N^{2}(v_{i+1})$. Since edges of $\Delta_{i+1}$ are of the form $N[x]$ for some $x\in X$, no two distinct $y,y^{\prime}$ can have private edges in $\Delta_{i+1}$ simultaneously. Hence, $|T\cap N^{2}(v_{i+1})|\leq 1$ as claimed. ∎
Lemma 4.
Let $T$ be a minimal transversal of $\Delta_{i+1}$. Then exactly one of the following holds:
  1. 1.
    $T=\{v_{i+1}\}$;
  2. 2.
    $T\subseteq B$ in which case $T=B$;
  3. 3.
    $T\subseteq R$ in which case $|T|\leq 2$; or
  4. 4.
    $T=\{r\}\cup(B\setminus N(r))$ for some $r\in N^{2}(v_{i+1})$.
Proof.
First, note that all the cases are mutually distinct. Hence, we may consider some $T\in Tr(\Delta_{i+1})$, assume that $T\neq\{v_{i+1}\}$, and prove the conclusion for each of Items 24 Note that each $x\in X$ hits a single edge $N[x]$ in $\Delta_{i+1}$, as $X$ is an independent set. Hence, if $T\subseteq B$ then $T=B$ concluding Item 2. If $T\subseteq R$ then we have that $|T|\leq 2$ by Proposition 9 deriving Item 3. We are thus left with proving Item 4. Suppose that $T\cap B\neq\emptyset$ and $T\cap R\neq\emptyset$. Note that if $x\in X\cap R$, then it may be incident to at most one edge in $\Delta_{i+1}$, namely $N[v_{i+1}]$. However, since $T\cap B\neq\emptyset$ by assumption, and since each element in $B$ is adjacent to $v_{i+1}$, the edge $N[v_{i+1}]$ may not be the private edge of such an $x$ in $T$. Hence $T\cap X\cap R=\emptyset$, that is, $T\cap R\subseteq Y$. By Proposition 9 there is exactly one vertex, call it $y$, belonging to $T\cap Y$. Finally, note that in addition to $N[v_{i+1}]$, each $b\in B$ hits a single edge $N[b]\in\Delta_{i+1}$ that only $b$ can hit in $X$. Hence $B\setminus N(r)\subseteq T$ and Item 4 follows. ∎
Note that the conditions of the items in Lemma 4 can be checked in polynomial time, and provide an upper bound of $O(n^{2})$ on the number of minimal transversals of $\Delta_{i+1}$. Clearly these minimal transversals can be constructed within the same time and using polynomial space. We formalize this as follows.
Corollary 2.
The set $Tr(\Delta_{i+1})$ can be listed in polynomial time and space.
We conclude the following by combining Corollaries 1 and 2. This answers positively Question 1.
Theorem 2.
There is a polynomial delay and space algorithm enumerating minimal dominating sets in chordal bipartite graphs.
Let us now argue that with small modifications we can get a polynomial-delay algorithm for TDom-Enum on the same class. Note that the minimal extension problem is almost identical, with the caveat that the hyperedges we must hit are now of the form $N(x)$ for $x=v_{i+1}$ or $x\in X$, i.e., we are now dealing with open instead of closed neighborhoods. A similar analysis as conducted in the proof of Lemma 4 shows that minimal extensions are red subsets dominating blue elements plus $v_{i+1}$ if $N(v_{i+1})\subseteq V_{i+1}$. We note that this becomes an instance of proper red-blue domination (where only red elements may be picked) which is precisely what is solved in [GHK+16a] for chordal bipartite graphs. However in our case, due to the instance being a bipartite chain graph, the problem is trivial. Indeed, it may be seen that solutions have size at most two, intersecting $X$ and $Y$ on at most one vertex. As such, we have an alternative proof of Question 2, which was already known by [GHK+16a].
[…]
5 Minimal connected dominating sets
Recall that CDom-Enum amounts to list the minimal transversals of the (implicit) hypergraph of minimal separators of a given graph $G$, i.e., Proposition 3. By Proposition 6, this hypergraph can be constructed in polynomial time if $G$ is chordal bipartite. In this section, we moreover show that this hypergraph has bounded conformality in that case. As a corollary, we can use the algorithm of Khachiyan et al. in [KBEG07] to list minimal connected dominating sets in incremental-polynomial time in chordal bipartite graphs, providing a positive answer to Question 2.
Let us define the notion of conformality introduced by Berge in [Ber84]. We say that a hypergraph $\mathcal{H}$ is of conformality $c$ if the following property holds for every subset $X\subseteq V(\mathcal{H})$: $X$ is contained (as a subset) in a hyperedge of $\mathcal{H}$ whenever each subset of $X$ of cardinality at most $c$ is contained (as a subset) in a hyperedge of $\mathcal{H}$. See also [Ber84, KBEG07] for other equivalent characterizations of bounded conformality. We recall that a hypergraph $\mathcal{H}$ is called Sperner if $E_{1}\not\subseteq E_{2}$ for any two distinct hyperedges $E_{1},E_{2}$ in $\mathcal{H}$.
Khachiyan, Boros, Elbassioni, and Gurvich proved the following.
Theorem 3 ([KBEG07]).
The minimal transversals of a Sperner hypergraph of bounded conformality can be enumerated in incremental-polynomial time.
We show that the family of minimal separators has bounded conformality in chordal bipartite graphs.
Lemma 5.
If $G$ is chordal bipartite then $\mathcal{S}(G)$ has conformality $5$.
Proof.
Let us put $\mathcal{H}:=\mathcal{S}(G)$. Assume that $\mathcal{H}$ is not of conformality 55{5}5, i.e., there is a subset $X\subseteq V(\mathcal{H})$ that is not contained in a hyperedge of $\mathcal{H}$, but each of its subsets $Y\subseteq X$ of size at most $5$ are contained in a hyperedge of $\mathcal{H}$. Here, by contained we mean as a subset. We consider $X=\{x_{1},\dots,x_{p}\}$ of minimum cardinality. Then ${p\geq 6}$ as otherwise, by assumption, $X$ would be contained in a hyperedge of $\mathcal{H}$. Moreover, for every $x_{i}\in X$, note that there must exist a hyperedge $E_{i}$ of $\mathcal{H}$ such that $E_{i}\cap X=X\setminus\{x_{i}\}$: we prove this depending on the value of $p$. In the first case, if $p\geq 7$, the nonexistence of such $E_{i}$ implies that $X^{\prime}=X\setminus\{x_{i}\}$ is not contained in a hyperedge of $\mathcal{H}$, and still every subset $Y\subseteq X^{\prime}$ of size at most $5$ is, contradicting the minimality of $X$; if, however, $p=6$ by definition we have that every $Y\subset X$ of size 5 is contained in some $E_{Y}\in E(\mathcal{H})$, but $X\nsubseteq E_{Y}$ since $\mathcal{H}$ is not of conformality 5.
Since $X$ has cardinality at least $6$, there exist at least three vertices $a_{1},a_{2},a_{3}$ in $X$ lying in one side of the bipartition of $G$, call it $A$. Let $E_{1}$ be an edge of $\mathcal{H}$, i.e., a minimal separator of $G$ that does not contain $a_{1}$, which exists by our previous arguments. By the minimality of $E_{1}$ and Proposition 7 there are at least two components that are close to $E_{1}$. Let $C_{1}$ be one that does not contain $a_{1}$. By Proposition 8 there exists $c_{1}$ in $C_{1}$ that is adjacent to both $a_{2}$ and $a_{3}$. However note that $c_{1}$ is not adjacent to $a_{1}$ as we chose $C_{1}$ to be the component not containing $a_{1}$.
Consider now the separator $E_{2}$. Again, by minimality there are at least two components that are close to $E_{2}$. Let $C_{2}$ be the one that does not contain $a_{2}$. By Proposition 8 there exists $c_{2}$ in $C_{2}$ that is adjacent to both $a_{1}$ and $a_{3}$. However $a_{2}$ and $c_{2}$ are non adjacent.
We repeat this procedure for $E_{3}$ and obtain $c_{3}$ that is adjacent to both $a_{1}$ and $a_{2}$ but not $a_{3}$. Note that as $a_{1},a_{2},a_{3}$ lie in $A$, $c_{1},c_{2},c_{3}$ lie in the other side of the bipartition and are thus pairwise non adjacent. We obtain an induced cycle $(a_{1},c_{1},a_{2},c_{2},a_{3},c_{3},a_{1})$ contradicting the fact that $G$ is $C_{\geq 6}$-free. ∎
We derive the following as a corollary of Theorem 3 and Lemma 5.
Theorem 4.
The minimal connected dominating sets of a chordal bipartite graph can be listed in incremental polynomial time.
6 Limitations of the sequential method
It is already known that the sequential method may not directly provide output-polynomial time algorithms for Trans-Enum in general; see e.g. [BDM24] for explicit statements. We adapt the construction in [BDM24] to show that this still holds for hypergraphs of (open or closed) neighborhoods of bipartite graphs, i.e., that we may not expect to get tractable algorithms for TDom-Enum and Dom-Enum using this technique.
We begin with Dom-Enum and later detail the modifications to be performed for the statement to hold for TDom-Enum. In the following by $\mathcal{N}(G)$ we mean the hypergraph of closed neighborhoods of $G$, whose minimal transversals are precisely the minimal dominating sets of $G$; see Section 2.
Recall that in Multicolored Independent Set problem, we are given a graph $G$, an integer $k$, and a partition $\mathcal{V}=(V_{1},V_{2},\dots,V_{k})$ of $V(G)$ such that $V_{i}$ is an independent set for every $1\leq i\leq k$. Sets $V_{1},\dots,V_{k}$ are usually referred to as color classes. The goal is then to decide whether there exists an independent set in $G$ containing precisely one vertex from each $V_{i}$, i.e., that is multicolored. See [CFK+15] for more details on this problem.
Figure 2: The construction of $H$, where vertices are ordered arbitrarily except for $u^{\star}$ and $v^{\star}$ which are put last, in this order. Color classes $V_{1},\dots,V_{k}$ induce independent sets in $H$. The dashed edge denote an edge of $G$ that is not part of $H$. Black vertices represent $T^{\star}$. Note that there is one copy of the depicted gadgets for each color class, vertex, and edge of $G$.
Lemma 6.
Let $(G,\mathcal{V})$ be an instance of Multicolored Independent Set. Then there exists a bipartite graph $H$, an ordering $v_{1},\dots,v_{n}$ of its vertices, and a minimal transversal $T^{\star}$ of $\mathcal{N}(H)_{n-2}$, such that the sets in $\operatorname{{\sf children}}(T^{\star},n-2)$ are all multicolored independent sets of $G$ except for one.
Proof.
Let $(G,\mathcal{V})$ be an instance of Multicolored Independent Set where $\mathcal{V}=(V_{1},\dots,V_{k})$ is a partition of $G$ and each $V_{i}$ is one of $k$ color classes, $n=|V(G)|$ and $m=|E(G)|$. We describe the construction of a bipartite graph $H$ from $G$ as follows, and create a set $T^{\star}$ of vertices along the way. See Figure 2 for an illustration.
We start by adding to $H$ the set of vertices $V(G)$, but none of the edges of $G$. For each color class $V_{i}$, $1\leq i\leq k$ we create a special vertex $c_{i}$ that we make adjacent to all vertices of $V_{i}$. Then, again for each color class, we create a path on four vertices $x_{1},x_{2},x_{3},x_{4}$ and attach it to $c_{i}$ by adding the edge $c_{i}x_{4}$. We add $x_{2}$ and $x_{3}$ to $T^{\star}$. The role of this gadget is to leave $c_{j}$ non-dominated by $T^{\star}$, yet to forbid $c_{j}$ in a minimal dominating set containing $T^{\star}$. In the following, we define $C=\{c_{i}:i\in\{1,\dots,k\}\}$.
For every edge $uv$ in $G$ we construct a subdivided claw $y_{1},y_{2},y_{3},y_{4},u,v$, with $N(y_{2})=\{y_{1},y_{3}\}$, $N(y_{4})=\{y_{3},u\}$, $N(y_{5})=\{y_{3},v\}$ and no other edges touch the $y$ vertices. Also, add $y_{2}$ and $y_{3}$ to $T^{\star}$. The objective of this gadget is to forbid the selection of both endpoints $u$ and $v$ of an edge in a minimal dominating set containing $T^{\star}$. Note that we construct one such gadget per edge of $G$.
For every vertex $v$ of $G$, we create a distinct path on four vertices $z_{1},z_{2},z_{3},z_{4}$, and attach it to $v$ with the edge $vz_{3}$. We add $z_{2}$ and $z_{3}$ to $T^{\star}$. This ensures that each such $v$ is dominated by $T^{\star}$. Moreover, note that picking $v$ in a minimal dominating set containing $T^{\star}$ will not steal private neighbors of an element in $T^{\star}$, and will dominate $c_{i}$ for $i$ such that $v\in V_{i}$.
Finally, we create two additional vertices $u^{\star}$ and $v^{\star}$ with $u^{\star}$ adjacent to each $c_{i}$, $1\leq i\leq k$, and $v^{\star}$ only adjacent to $u^{\star}$. We fix the ordering of vertices of $H$ to be arbitrary but ending with $u^{\star},v^{\star}$ in this order.
This completes the construction of $H$. It has $O(n+m)$ vertices and edges. Moreover, it is easily seen that $H$ is bipartite, by taking $V(G)$ on one side of the bipartition, and noting that except for $u^{\star}$ and $v^{\star}$ the gadgets only create even paths between pairs of vertices attached to $V(G)$. Then $u^{\star}$ is made adjacent to the $c_{j}$’s which lie in the opposite partition of $V(G)$, and $v^{\star}$ is a pendant vertex. Thus they do not create odd cycles. See also Figure 2 for a representation of $H$ in layers where each edge is between consecutive layers.
Let us define $\mathcal{H}:=\mathcal{N}(H)$ and show that $T^{\star}$ is a minimal transversal of $\mathcal{H}_{n-2}=\mathcal{H}\left[V(H)\setminus\{u^{\star},v^{\star}\}\right]$. First note that all neighborhoods except for those of $c_{i}$’s, $u^{\star}$ and $v^{\star}$ are included in the $n-2$ first vertices, hence belong to $\mathcal{H}_{n-2}$. Moreover, all these neighborhoods are intersected by $T^{\star}$ as every gadget is dominated by $T^{\star}$ by construction, and each vertex $v\in V(G)$ is dominated by the vertex $z_{3}$ of its corresponding vertex gadget. Furthermore, $T^{\star}$ is minimal with that property as each of its element $t$ has a private neighbor in its corresponding gadget, whose edge is thus only intersected by $t$ in $T^{\star}$.
Let us now consider the different possibilities for a minimal transversal $T$ of $\mathcal{H}_{n-1}$ to contain $T^{\star}$ as a subset. Note that only the neighborhoods of $c_{i}$ for each $1\leq i\leq k$ are to be dominated. To do so, we either pick $u^{\star}$, or must select a subset $X$ of $V(G)\cup C$ in order to minimally hit $N[c_{i}]$ for all $1\leq i\leq k$. Note however that as $x_{4}$ is the only private neighbor of $x_{3}$ in the color class gadget of $V_{i}$, $X$ cannot contain $c_{i}$ nor $x_{1}$ for any color. Hence $X\subseteq V(G)$. Note also that no two distinct elements of $X$ intersect a same color class $V_{i}$ as their only private neighbor would be $c_{i}$, and hence one of the two would be redundant. Moreover by the edge gadget, since $y_{4}$ and $y_{5}$ are the only two private neighbors of $y_{3}$, $X$ cannot contain both $u$ and $v$ for any edge $uv$ of $G$. Thus $X$ intersects each color class precisely once, and does not contain two vertices $u,v$ with $uv\in E(G)$. We conclude that $X$ is a multicolored independent set of $G$, hence that $T^{\star}$ can be extended into a minimal transversal of $\mathcal{H}_{n-1}$ using a different set than $\{u^{*}\}$ if and only if $G$ contains a multicolored independent set.
Finally, recall that by Lemma 2 and Lemma 3 either $\operatorname{{\sf children}}(T^{\star},n-2)=\{T^{\star}\}$, or sets in $\operatorname{{\sf children}}(T^{\star},n-2)$ are of the form $T^{\star}\cup X$ for some $X\in Tr(\Delta_{n-1})$ where
$\Delta_{n-1}:=\{E\in\mathcal{H}_{n-1}:E\cap T^{\star}=\emptyset\}=\{N[c_{i}]:1\leq i\leq k\}.$
Since $T^{\star}$ is not a transversal of $\mathcal{H}_{n-1}$ we are in the second case. Moreover, by the above discussion, each such $X\in Tr(\Delta_{n-1})$ is either $\{u^{*}\}$ or is a multicolored independent set of $G$. We thus need to show that any multicolored independent set of $G$ yields a children to conclude the proof.
Let $X$ be a multicolored independent set of $G$ and suppose that $T^{\star}\cup X$ is not a child of $(T^{*},n-2)$. Then during the computation of the parent of $T^{\star}\cup X$, some element $v\in T^{*}$ is removed. Hence such a $v$ belongs to one of the edge, vertex, of color class gadgets. But for these element to lose a private neighbor it should be that $X$ contains two elements of a same color class, or two endpoints of an edge, a contradiction. ∎
A consequence of Lemma 6 is that the sequential approach may not be directly applied to solve Dom-Enum, even in bipartite graphs. Moreover, note thatin our construction, no vertex $v$ of $T^{\star}$ is made self-private, i.e., no $v$ has its closed neighborhood as a private edge. Furthermore, the elements in the color classes are dominated by $T^{\star}$, and their role is thus limited to dominate the $c_{i}$’s in an extension. Hence, by keeping the same construction, all the arguments of the above proof hold for the hypergraph of open neighborhoods of $G$. We derive that the sequential approach may not be directly be used for TDom-Enum neither , as summarized in Lemma 7, where $\mathcal{N}^{o}(H)$ is the hypergraph of open neighborhoods of $H$.
Lemma 7.
Let $(G,\mathcal{V})$ be an instance of Multicolored Independent Set. Then there exists a bipartite graph $H$, an ordering $v_{1},\dots,v_{n}$ of its vertices, and a minimal transversal $T^{\star}$ of $\mathcal{N}^{o}(H)_{n-2}$, such that the sets in $\operatorname{{\sf children}}(T^{\star},n-2)$ are all multicolored independent sets of $G$ except for one.
7 Trans-Enum and bipartite graphs
Our final result is a proof that Trans-Enum reduces to CDom-Enum on bipartite graphs, which we are not aware of being previously known or directly implied by other results. In fact, by Proposition 3, if we can construct a graph where (almost) every minimal separator is in a one to one correspondence to our hyperedges, we are done.
Let $\mathcal{H}$ be the input hypergraph to Trans-Enum with $E(\mathcal{H})=\{E_{1},\dots,E_{m}\}$ and $V(\mathcal{H})=\{u_{1},\dots,u_{n}\}$ and assume w.l.o.g. that $\mathcal{H}$ is a Sperner hypergraph. To obtain $G$, the input to CDom-Enum, we start with the incidence bipartite graph of $\mathcal{H}$, that is, $\{v_{1},\dots,v_{n}\}$, $\{e_{1},\dots,e_{m}\}\subset V(G)$, where $v_{i}$ corresponds to $u_{i}\in V(\mathcal{H})$ and $e_{j}$ corresponds to $E_{j}\in E(\mathcal{H})$, and $v_{i}e_{j}\in E(G)$ if and only if $u_{i}\in E_{j}$. To conclude the construction of $G$, we add two vertices $v^{\prime}$ and $v^{\star}$, the edge $v^{\prime}v^{\star}$, and the edges $v^{\star}v_{i}$ for all $1\leq i\leq n$.
Lemma 8.
The minimal separators of $G$ are $N(e_{1}),\dots,N(e_{m})$ and $\{v^{\star}\}$.
Proof.
To see that $\{v^{\star}\}\in\mathcal{S}(G)$, observe that $G\setminus\{v^{\star}\}$ has two connected components: $\{v^{\prime}\}$ and $V(G)\setminus\{v^{\prime},v^{\star}\}$. As such, no other $S\in\mathcal{S}(G)$ contains $v^{\star}$, so it holds that every $v_{i}\notin S$ is in the same connected component as $v^{\star}$ and $v^{\prime}$. Let $E_{j}\in E(\mathcal{H})$ and observe that $N(e_{j})$ separates $e_{j}$ from the rest of $G\setminus N(e_{j})$; moreover, note that, since $\mathcal{H}$ is a Sperner hypergraph, every other $e_{\ell}\in V(G)$ has a neighbor $v_{k}\notin N(e_{j})$ (as no two hyperedges satisfy a containment relation), so it holds that $G\setminus N(e_{j})$ has exactly two components: $\{e_{j}\}$ and the rest of the graph. Thus, if we add back one vertex from $N(e_{j})$, $e_{j}$ is no longer disconnected as it creates a $e_{j}$$v^{\star}$ path. Now, towards a contradiction, assume that there is another $S\in\mathcal{S}(G)$ that is different from every $N(e_{j})$ and from $\{v^{\star}\}$. Let $\{v^{\prime},v^{\star},v_{1},\dots,v_{n}\}\setminus S\subseteq C_{1}$ be a non-empty connected component of $G\setminus S$; note that, because $V(G)\setminus(S\cup C_{1})\subseteq\{e_{1},\dots,e_{m}\}$, the other connected components of $G\setminus S$ (which must exist as $S$ is a separator) must have size one, but this implies that each $e_{j}\notin C_{1}\cup S$ has $N(e_{j})=N(e_{j})\subseteq S$, a contradiction. ∎
Theorem 5.
CDom-Enum on bipartite graphs is at least as hard as Trans-Enum.
Proof.
By our construction and Lemma 8, each $T_{G}\in Tr(\mathcal{S}(G))$ contains $v^{\star}$ and $T_{G}\setminus\{v^{\star}\}$ minimally hits each $N(e_{j})=\{v_{i}:u_{i}\in E_{j}\}$; thus $T_{\mathcal{H}}(T_{G})=\{u_{i}:v_{i}\in T_{G}\setminus\{v^{\star}\}\}$ is a bijection between $Tr(\mathcal{S}(G))$ and $Tr(\mathcal{H})$. By Proposition 3, $\operatorname{MCDS}(G)=Tr(\mathcal{S}(G))$, so there is a bijection between $\operatorname{MCDS}(G)$ and $Tr(\mathcal{H})$ and the result follows. ∎
8 Discussion
In this paper, we investigated the complexity of Dom-Enum and its two variants TDom-Enum and CDom-Enum in chordal bipartite graphs. We obtained polynomial-delay algorithms for the first two problems, and an incremental-polynomial time algorithm for the last problem. Then we gave evidence that the techniques used for Dom-Enum and TDom-Enum may not be pushed directly to bipartite graphs. As of CDom-Enum, we proved a stronger statement that CDom-Enum is as hard as Trans-Enum in bipartite graphs, with the later problem being open since more than forty years [EG95].
We leave open the existence of polynomial-delay algorithms for Dom-Enum and TDom-Enum in bipartite graphs, and for CDom-Enum in chordal bipartite graphs. Concerning this last case, we note that obtaining polynomial delay is open for Trans-Enum on hypergraph of bounded dimension, which define a proper subclass of hypergraphs of bounded conformality. Hence, in order to improve Theorem 4, one either has to find another strategy, or should aim at improving the results in [EG95, KBEG07] first.